OpenStudy (anonymous):

write an equation and solve to find three consecutive integers that have a sum of - 45

OpenStudy (anonymous):

x+(x+1)+(x+2)=-45 3x+3=-45 x+1=-15 x=-16 The three integers are -16,-15,-14